Assume that a company sells customized sweatshirts for $15 per unit. It pays a sales commission of $5 per unit sold. The company must buy sweatshirts from its supplier in batches of 100 units at an average unit cost of $7 per sweatshirt. The number of sweatshirts the company would need to sell to earn a target profit of $1,650 is closest to:
Answer #1

Selling price per unit = $15

Sales commission per unit = $5

Purchase price per unit = $7

Total cost per unit = Sales commission per unit + Purchase price per unit

= 5+7

= $12

Profit per unit = Selling price per unit- Total cost per unit

= 15-12

= $3

Target profit = $1,650

Number of units sold = Target profit / Profit per unit

= 1,650/3

= 550

The number of sweatshirts the company would need to sell to earn a target profit of $1,650 = 550

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks